A satellite of mass 5550 kg orbits the Earth and has a period of 6610 s .

Determine the radius of its circular orbit.

Express your answer using three significant figures and include the appropriate units.

Determine the magnitude of the Earth's gravitational force on the satellite.

Express your answer using three significant figures and include the appropriate units.

Determine the altitude of the satellite.

Express your answer using three significant figures and include the appropriate units.
Answer #1

A)

Radius, r = [GM T^2/4 pi^2]^(1/3)

r = [(6.67 x 10^-11)(6 x 10^24)(6610^2)/4 pi^2]^(1/3)

r = 7.62 x 10^6 m

B)

Force, F = GMm/r^2

F = (6.67 x 10^-11 x 6 x 10^24 x 5550)/(7.62 x 10^6)^2

F = 3.82 x 10^4 N

C)

Altitude, h = r - R = (7.62 x 10^6) - (6.38 x 10^6)

h = 1.24 x 10^6 m

Comment in case any doubt please rate my answer...